100 school children: On first day of term 8 have a cold. During the first week (the study period) a further 15 develop a cold.

1. What is the prevalence on day1 ?

2. What is the cumulative incidence? First work out the population at risk on day 1
Answer #1

The total number of childrens in the school - 100

On first day of the term 8 children had cold

During the first week -15 more children developed cold

1. Prevalence rate is the proportion of persons in a population who have a particular disease or attribute at a specified point in time or over a specified period of time.

Prevalence rate of day 1 = Number of all childrens with cold / Total number of children at that specified point in time

= 8 / 100 = 0.08 %

2. Cumulative incidence - Number of new cases of disease during a given period of time / total population at risk

= Number of new cases with cold - 15

Total number of children = 100

Cumulative incidence = 15/ 100

= 0.15